Suppose Abby tracks her travel time to work for 60 days and determines that her mean travel time, in minutes, is 35.6. Assume the travel times are normally distributed with a standard deviation of 10.3 min. Determine the travel time ? such that 22.96% of the 60 days have a travel time that is at least x.

x=_____min
Answer #1

Solution :

Given that,  

mean = \mu = 35.6

standard deviation = \sigma = 10.3

Using standard normal table ,

P(Z > z) = 22.96 %

1 - P(Z < z) = 0.2296

P(Z < z) = 1 - 0.2296

P(Z < 0.74) = 0.7704

z = 0.74

Using z-score formula,

x = z * \sigma + \mu

x = 0.74 * 10.3 + 35.6 = 43.22

x = 43.22 mim